OpenStudy (anonymous):

How can i completly factor this equation? 3x^2 - 3x - 60

OpenStudy (anonymous):

first take out the GCF for the trinomial...3(x^2 - x -20). it should be easier to factor now...

OpenStudy (anonymous):

how do I factor it completly...

OpenStudy (anonymous):

3(x+4)(x-5)

OpenStudy (anonymous):

can u show me step by step?

OpenStudy (anonymous):

sure....

OpenStudy (amistre64):

thats never an easy task is it

OpenStudy (amistre64):

assume it factors: (x+a)(x+b) = x^2 -x -20 expand the left side x^2+(a+b)x +ab = x^2 -x -20 and compare coeeficients is prolly the simplest route

OpenStudy (anonymous):